# Estimating the FDR of variable selection

### v2

@@ Appendix D Deferred proofs
$\frac{1}{2}\,\mathbb{E}_{\textnormal{RSS}}\left[\left(\frac{T}{R}-\frac{T^{\prime}}{R^{\prime}}\right)^{2}\mid R,R^{\prime}>Kq\right]\cdot\mathbb{P}_{\textnormal{RSS}}(R,R^{\prime}>Kq)$
$=$ $\frac{1}{2}\,\mathbb{E}_{\textnormal{RSS}}\left[\left(\frac{TR^{\prime}-T^{\prime}R}{RR^{\prime}}\right)^{2}\mid R,R^{\prime}>Kq\right]\cdot\mathbb{P}_{\textnormal{RSS}}(R,R^{\prime}>Kq)$
$\leq$ $\frac{1}{2K^{4}q^{4}}\cdot\mathbb{E}_{\textnormal{RSS}}\left[\left(TR^{\prime}-T^{\prime}R\right)^{2}\mid R,R^{\prime}>Kq\right]\cdot\mathbb{P}_{\textnormal{RSS}}(R,R^{\prime}>Kq)$
$\leq$ $\frac{1}{2K^{4}q^{4}}\cdot\mathbb{E}_{\textnormal{RSS}}\left[\left(TR^{\prime}-T^{\prime}R\right)^{2}\right]$
$=$ $\frac{1}{2K^{4}q^{4}}\cdot\left(2\textnormal{Var}_{\textnormal{RSS}}\left[TR^{\prime}\right]-2\left(\mathbb{E}_{\textnormal{RSS}}^{2}\left[TR\right]-\mathbb{E}_{\textnormal{RSS}}^{2}\left[T\right]\mathbb{E}_{\textnormal{RSS}}^{2}\left[R\right]\right)\right)$
$\leq$ $\frac{1}{K^{4}q^{4}}\cdot\left(\textnormal{Var}_{\textnormal{RSS}}\left[TR^{\prime}\right]+2\mathbb{E}_{\textnormal{RSS}}\left[T\right]\mathbb{E}_{\textnormal{RSS}}\left[R\right]\cdot|\text{Cov}_{\textnormal{RSS}}(T,R)|\right)$
@@
Note $T_{k}\in[0,c_{\phi}b]$, we have $\mathbb{E}_{\textnormal{RSS}}[T]\leq Kc_{\phi}b$ and $\textnormal{Var}_{\textnormal{RSS}}[T]=\sum_{k}\textnormal{Var}_{\textnormal{RSS}}[T_{k}]\leq Kc_{\phi}^{2}b^{2}$. Similarly, $\mathbb{E}_{\textnormal{RSS}}[R^{\prime}]\leq Kb$ and $\textnormal{Var}_{\textnormal{RSS}}[R^{\prime}]\leq Kb^{2}$. Thus
$\textnormal{Var}_{\textnormal{RSS}}\left[TR^{\prime}\right]\leq K^{2}c_{\phi}^{2}b^{4}+K^{3}c_{\phi}^{2}b^{4}=O(b^{4}K^{3})$
@@
$|\text{Cov}_{\textnormal{RSS}}(T,R)|\leq\sqrt{\textnormal{Var}_{\textnormal{RSS}}[T]\textnormal{Var}_{\textnormal{RSS}}[R]}\leq Kb^{2}c_{\phi}.$
Therefore the first term is bounded by $O(b^{4}/K)$. For the second term,
$\frac{1}{2}\,\mathbb{E}_{\textnormal{RSS}}\left[\left(\frac{T}{R}-\frac{T^{\prime}}{R^{\prime}}\right)^{2}\mid R\wedge R^{\prime}\leq Kq\right]$
$\leq$ $\frac{1}{2}\,\mathbb{E}_{\textnormal{RSS}}\left[\left(\frac{T}{R}+\frac{T^{\prime}}{R^{\prime}}\right)^{2}\mid R\wedge R^{\prime}\leq Kq\right]$
$\leq$ $\frac{1}{2}\,\mathbb{E}_{\textnormal{RSS}}\left[\left(T+T^{\prime}\right)^{2}\mid R\wedge R^{\prime}\leq Kq\right]$
$\leq$ $2K^{2}c_{\phi}^{2}b^{2}.$
The last inequality is because $T\leq Kc_{\phi}b$. By Proposition 3.1, we have
$\mathbb{P}_{\textnormal{RSS}}(R\wedge R^{\prime}\leq Kq)\leq 2\mathbb{P}_{\textnormal{RSS}}(R\leq Kq)\leq 2\rho.$
@@
$\textnormal{Var}_{\textnormal{RSS}}\left[\frac{T}{R}\right]\leq c_{\phi}^{2}\cdot\left[\frac{K^{2}b^{4}+5K^{3}b^{4}}{K^{4}q^{4}}+2K^{2}b^{2}\cdot 2\rho\right].$
The bound is of $O\left({b^{4}}/{K}\right)$ if $\rho=O(K^{-3})$.
Step 2. $\textnormal{Var}\left[\mathbb{E}\left({\sum_{k}T_{k}}/{\sum_{k}R_{k}}\mid\textnormal{RSS}\right)\right]=O(1/(n-d))$.
We start by bounding the perturbation of $\mu_{j}(j\in\mathcal{R}_{k})$ when we perturb RSS while holding $\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$ fixed. We write
[…]
where $r_{j}=\sqrt{\textnormal{RSS}+u_{j}^{2}}$, $I_{j}(v_{j};X_{B_{k}-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y})=\mathbf{1}\left\{j\in\mathcal{R}_{k}\right\}$, and $q_{j}=d\mu_{j}/dv_{j}$ is the probability density. Direct calculation shows
@@
$\partial_{\textnormal{RSS}}\,\mu_{j}(j\in\mathcal{R}_{k})=\int_{-r_{j}}^{r_{j}}I_{j}(v_{j};X_{B_{k}-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y})\cdot\partial_{\textnormal{RSS}}\,q_{j}(v_{j})\,dv_{j},$
@@
[…]
Since $I_{j}(v_{j};X_{B_{k}-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y})\in\left\{0,1\right\}$, we have
$|\partial_{\textnormal{RSS}}\,\mu_{j}(j\in\mathcal{R}_{k})|$ $\leq\max\left\{\int_{r_{j}^{2}\geq v_{j}^{2}\geq\frac{r_{j}^{2}}{n-d-1}}\partial_{\textnormal{RSS}}\,q_{j}\,dv_{j},\;-\int_{v_{j}^{2}\leq\frac{r_{j}^{2}}{n-d-1}}\partial_{\textnormal{RSS}}\,q_{j}\,dv_{j}\right\}$
@@
where $F_{t_{n-d}}$ is the cumulative distribution function of the t-distribution of degree of freedom $n-d$. Then
@@
Now consider the event $A_{\textnormal{RSS}}=\left\{0.5\,(n-d)\leq\textnormal{RSS}\leq 1.5\,(n-d)\right\}$. Recall we assume $\sigma^{2}=1$. Since $\textnormal{RSS}\sim\chi^{2}_{n-d}$ is sub-exponential, we have $\mathbb{P}(A^{\mathsf{c}})\leq 2\exp(-(n-d)/32)$ by the Bernstein concentration equalities.
@@
$|\partial_{\textnormal{RSS}}\,\mu_{j}(j\in\mathcal{R}_{k})|\leq\frac{2}{n-d}$
@@
Note both $\mu_{j}(j\in\mathcal{R}_{k})$ and $\phi_{j}(p_{j})$ are bounded. Therefore, $\mu_{j}(j\in\mathcal{R}_{k})\cdot\phi_{j}(p_{j})$ is Lipschitz in RSS with a Lipschitz constant
$\left(2c_{\phi}+L_{\psi}\right)\cdot\frac{1}{n-d}.$
Moreover, $\sum_{k}T_{k}=\sum_{j=1}^{Kb}\mu_{j}(j\in\mathcal{R}_{k})\cdot\phi_{j}(p_{j})$ is then Lipschitz in RSS with a Lipschitz constant $Kb\left(2c_{\phi}+L_{\psi}\right)/(n-d)$. A key observation is that ${\sum_{k}T_{k}}/{R}$ is Lipschitz in RSS since $R$ is not a function of RSS.
Now consider the event $A_{R}=\left\{R>Kq\right\}$. Since RSS is independent of $\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$, $A_{\textnormal{RSS}}$ and $A_{R}$ are independent. In addition, we have $\mathbb{P}(A_{R}^{\mathsf{c}})\leq\rho$.
On $A_{\textnormal{RSS}}\cap A_{R}$, we have ${\sum_{k}T_{k}}/{R}$ is $L$-Lipschitz in RSS with
$L=\left(\frac{2bc_{\phi}+bL_{\psi}}{q}\right)\cdot\frac{1}{n-d}.$
@@
$\mathbb{E}\left[\frac{\sum_{k}T_{k}}{R}\mid\textnormal{RSS},A_{R}\right]$
@@
$E_{R}(\textnormal{RSS})=\mathbb{E}\left[\frac{\sum_{k}T_{k}}{R}\mid\textnormal{RSS},A_{R}\right],\quad E_{R^{\mathsf{c}}}(\textnormal{RSS})=\mathbb{E}\left[\frac{\sum_{k}T_{k}}{R}\mid\textnormal{RSS},A_{R}^{\mathsf{c}}\right],$
@@
$E(\textnormal{RSS})=\mathbb{E}\left[\frac{\sum_{k}T_{k}}{R}\mid\textnormal{RSS}\right]=E_{R}(\textnormal{RSS})\cdot\mathbb{P}(A_{R})+E_{R^{\mathsf{c}}}(\textnormal{RSS})\cdot\mathbb{P}(A_{R}^{\mathsf{c}}).$
Note all $E_{R}(\textnormal{RSS})$, $E_{R^{\mathsf{c}}}(\textnormal{RSS})$, and $E(\textnormal{RSS})$ are bounded by $Kc_{\phi}b$.
@@
$=$ $(E_{R}(\textnormal{RSS})-E_{R}(\textnormal{RSS}^{\prime}))^{2}+O(c_{\phi}^{2}b^{2}\cdot K^{2}\rho)$
@@
$\quad+O(c_{\phi}^{2}b^{2}\cdot K^{2}\rho)+O\left(c_{\phi}^{2}b^{2}\cdot K^{2}e^{-{(n-d)}/{32}}\right)$
@@
$=$ $\left(\frac{2bc_{\phi}+bL_{\psi}}{q}\right)^{2}\cdot\frac{2}{n-d},$
where the last line uses the fact that $(n-d)\textnormal{RSS}\sim\chi^{2}(n-d)$ and hence $\textnormal{Var}(\textnormal{RSS})=2/(n-d)$. Then we have
$\textnormal{Var}\left[E(\textnormal{RSS})\right]\leq\left(\frac{2bc_{\phi}+bL_{\psi}}{q}\right)^{2}\cdot\frac{2}{n-d}+O(c_{\phi}^{2}b^{2}\cdot K^{2}\rho)+O\left(c_{\phi}^{2}b^{2}\cdot K^{2}e^{-(n-d)/32}\right).$
@@
$\textnormal{Var}\left[\mathbb{E}\left(\frac{\sum_{k}T_{k}}{R}\mid\textnormal{RSS}\right)\right]\leq\left(\frac{2bc_{\phi}+bL_{\psi}}{q}\right)^{2}\cdot\frac{2}{n-d}+O(c_{\phi}^{2}b^{2}\cdot K^{2}(\rho\vee e^{-(n-d)/32})).$
@@
$\textnormal{Var}\left[\frac{\sum_{k}T_{k}}{R}\right]$ $=\mathbb{E}\left[\textnormal{Var}_{\textnormal{RSS}}\left(\frac{T}{R}\right)\right]+\textnormal{Var}\left[\mathbb{E}\left(\frac{\sum_{k}T_{k}}{R}\mid\textnormal{RSS}\right)\right]$
$\leq O\left(\frac{c_{\phi}^{2}b^{4}}{Kq^{4}}\right)+\left(\frac{2bc_{\phi}+bL_{\psi}}{q}\right)^{2}\cdot\frac{2}{n-d}+O(c_{\phi}^{2}b^{2}\cdot K^{2}(\rho\vee e^{-(n-d)/32}))$
@@
$\textnormal{Var}\left[\widehat{\textnormal{FDR}}\right]=$ $\textnormal{Var}\left[\frac{\sum_{k}T_{k}}{\sum_{k}R_{k}}+O\left(\frac{Kc_{\phi}b^{3}}{R^{2}}\right)\right]$
$\leq$ $2\textnormal{Var}\left[\frac{\sum_{k}T_{k}}{\sum_{k}R_{k}}\right]+O\left(2\mathbb{E}\left[\frac{K^{2}c_{\phi}^{2}b^{6}}{R^{4}}\right]\right)$
$\leq$ $2\textnormal{Var}\left[\frac{\sum_{k}T_{k}}{\sum_{k}R_{k}}\right]+O\left(\frac{c_{\phi}^{2}b^{6}}{K^{2}q^{4}}\right)+O(K^{2}c_{\phi}^{2}b^{4}\rho)$
$\leq$ $O\left(c_{\phi}^{2}\left(\frac{b^{4}}{q^{4}}+\frac{b^{6}}{Kq^{4}}\right)\cdot\frac{1}{K}\right)+\left(\frac{2bc_{\phi}+bL_{\psi}}{q}\right)^{2}\cdot\frac{4}{n-d}$
$+O(c_{\phi}^{2}b^{4}\cdot K^{2}(\rho\vee e^{-(n-d)/32})).$
@@
Result (1): The setting satisfies the conditions in Part 2 of Proposition 3.1 with $\xi=K^{-\varepsilon}$ and $q=\xi/3=O(K^{-\varepsilon})$. Then $\rho=\exp\left(-K^{1-\varepsilon}/36\right)$. Plugging them into the bound in Theorem 3.1, we have
$\textnormal{Var}(\widehat{\textnormal{FDR}})=O\left(K^{-1+4\varepsilon}+\frac{K^{2\varepsilon}}{n-d}+K^{2}\exp\left(-K^{1-\varepsilon}/36\right)\right)=O(K^{-1+4\varepsilon}).$

### v4

@@ Appendix D Variance of FDR estimation
$\frac{1}{2}\,\mathbb{E}_{\textnormal{RSS}}\left[\left(\frac{T}{R}-\frac{T^{\prime}}{R^{\prime}}\right)^{2}\mid R,R^{\prime}>Bq\right]\cdot\mathbb{P}_{\textnormal{RSS}}(R,R^{\prime}>Bq)$
$=$ $\frac{1}{2}\,\mathbb{E}_{\textnormal{RSS}}\left[\left(\frac{TR^{\prime}-T^{\prime}R}{RR^{\prime}}\right)^{2}\mid R,R^{\prime}>Bq\right]\cdot\mathbb{P}_{\textnormal{RSS}}(R,R^{\prime}>Bq)$
$\leq$ $\frac{1}{2B^{4}q^{4}}\cdot\mathbb{E}_{\textnormal{RSS}}\left[\left(TR^{\prime}-T^{\prime}R\right)^{2}\mid R,R^{\prime}>Bq\right]\cdot\mathbb{P}_{\textnormal{RSS}}(R,R^{\prime}>Bq)$
$\leq$ $\frac{1}{2B^{4}q^{4}}\cdot\mathbb{E}_{\textnormal{RSS}}\left[\left(TR^{\prime}-T^{\prime}R\right)^{2}\right]$
$=$ $\frac{1}{2B^{4}q^{4}}\cdot\left(2\textnormal{Var}_{\textnormal{RSS}}\left[TR^{\prime}\right]-2\left(\mathbb{E}_{\textnormal{RSS}}^{2}\left[TR\right]-\mathbb{E}_{\textnormal{RSS}}^{2}\left[T\right]\mathbb{E}_{\textnormal{RSS}}^{2}\left[R\right]\right)\right)$
$\leq$ $\frac{1}{B^{4}q^{4}}\cdot\left(\textnormal{Var}_{\textnormal{RSS}}\left[TR^{\prime}\right]+2\mathbb{E}_{\textnormal{RSS}}\left[T\right]\mathbb{E}_{\textnormal{RSS}}\left[R\right]\cdot|\text{Cov}_{\textnormal{RSS}}(T,R)|\right)$
@@
Note $T_{b}\in[0,c_{\phi}m]$, we have $\mathbb{E}_{\textnormal{RSS}}[T]\leq Bc_{\phi}m$ and $\textnormal{Var}_{\textnormal{RSS}}[T]=\sum_{b}\textnormal{Var}_{\textnormal{RSS}}[T_{b}]\leq Bc_{\phi}^{2}m^{2}$. Similarly, $\mathbb{E}_{\textnormal{RSS}}[R^{\prime}]\leq Bm$ and $\textnormal{Var}_{\textnormal{RSS}}[R^{\prime}]\leq Bm^{2}$. Thus
$\textnormal{Var}_{\textnormal{RSS}}\left[TR^{\prime}\right]\leq B^{2}c_{\phi}^{2}m^{4}+B^{3}c_{\phi}^{2}m^{4}=O(m^{4}B^{3})$
@@
$|\text{Cov}_{\textnormal{RSS}}(T,R)|\leq\sqrt{\textnormal{Var}_{\textnormal{RSS}}[T]\textnormal{Var}_{\textnormal{RSS}}[R]}\leq Bm^{2}c_{\phi}.$
Therefore the first term is bounded by $O(m^{4}/B)$. For the second term,
$\frac{1}{2}\,\mathbb{E}_{\textnormal{RSS}}\left[\left(\frac{T}{R}-\frac{T^{\prime}}{R^{\prime}}\right)^{2}\mid R\wedge R^{\prime}\leq Bq\right]$
$\leq$ $\frac{1}{2}\,\mathbb{E}_{\textnormal{RSS}}\left[\left(\frac{T}{R}+\frac{T^{\prime}}{R^{\prime}}\right)^{2}\mid R\wedge R^{\prime}\leq Bq\right]$
$\leq$ $\frac{1}{2}\,\mathbb{E}_{\textnormal{RSS}}\left[\left(T+T^{\prime}\right)^{2}\mid R\wedge R^{\prime}\leq Bq\right]$
$\leq$ $2B^{2}c_{\phi}^{2}m^{2}.$
The last inequality is because $T\leq Bc_{\phi}m$. By Proposition 3.1, we have
$\mathbb{P}_{\textnormal{RSS}}(R\wedge R^{\prime}\leq Bq)\leq 2\mathbb{P}_{\textnormal{RSS}}(R\leq Bq)\leq 2\rho.$
@@
$\textnormal{Var}_{\textnormal{RSS}}\left[\frac{T}{R}\right]\leq c_{\phi}^{2}\cdot\left[\frac{B^{2}m^{4}+5B^{3}m^{4}}{B^{4}q^{4}}+2B^{2}m^{2}\cdot 2\rho\right].$
The bound is of $O\left({m^{4}}/{B}\right)$ if $\rho=O(B^{-3})$.
Step 2. $\textnormal{Var}\left[\mathbb{E}\left({\sum_{b}T_{b}}/{\sum_{b}R_{b}}\mid\textnormal{RSS}\right)\right]=O(1/(n-d))$.
We start by bounding the perturbation of $\mu_{j}(j\in\mathcal{R}_{b})$ when we perturb RSS while holding $\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$ fixed. We write
$\mu_{j}(j\in\mathcal{R}_{b})=\mathbb{P}_{H_{j}}(v_{j}\in A\mid\boldsymbol{X}_{(b);-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},r_{j}^{2})=\int_{-r_{j}}^{r_{j}}I_{j}(v_{j};\boldsymbol{X}_{(b);-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y})\cdot q_{j}(v_{j})\,dv_{j},$
where $r_{j}=\sqrt{\textnormal{RSS}+u_{j}^{2}}$, $I_{j}(v_{j};X_{(b);-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y})=\mathbf{1}\left\{j\in\mathcal{R}_{b}\right\}$, and $q_{j}=d\mu_{j}/dv_{j}$ is the probability density. Direct calculation shows
@@
$\partial_{\textnormal{RSS}}\,\mu_{j}(j\in\mathcal{R}_{b})=\int_{-r_{j}}^{r_{j}}I_{j}(v_{j};X_{(b);-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y})\cdot\partial_{\textnormal{RSS}}\,q_{j}(v_{j})\,dv_{j},$
@@
$\int_{-r_{j}}^{r_{j}}\partial_{\textnormal{RSS}}q_{j}dv_{j}=\partial_{\textnormal{RSS}}\int_{-r_{j}}^{r_{j}}q_{j}dv_{j}=\partial_{\textnormal{RSS}}1=0.$
Since $I_{j}(v_{j};X_{(b);-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y})\in\left\{0,1\right\}$, we have
$|\partial_{\textnormal{RSS}}\,\mu_{j}(j\in\mathcal{R}_{b})|$ $\leq\max\left\{\int_{r_{j}^{2}\geq v_{j}^{2}\geq\frac{r_{j}^{2}}{n-d-1}}\partial_{\textnormal{RSS}}\,q_{j}\,dv_{j},\;-\int_{v_{j}^{2}\leq\frac{r_{j}^{2}}{n-d-1}}\partial_{\textnormal{RSS}}\,q_{j}\,dv_{j}\right\}$
@@
where $F_{t_{n-d}}$ is the cumulative distribution function of the t-distribution of degree of freedom ndn-d. Then
@@
Now consider the event $A_{\textnormal{RSS}}=\left\{0.5\,(n-d)\leq\textnormal{RSS}\leq 1.5\,(n-d)\right\}$. Recall we assume $\sigma^{2}=1$. Since $\textnormal{RSS}\sim\chi^{2}_{n-d}$ is sub-exponential, we have $\mathbb{P}(A^{\mathsf{c}})\leq 2\exp(-(n-d)/32)$ by the Bernstein concentration inequalities.
@@
$|\partial_{\textnormal{RSS}}\,\mu_{j}(j\in\mathcal{R}_{b})|\leq\frac{2}{n-d}$
@@
Note both $\mu_{j}(j\in\mathcal{R}_{b})$ and $\phi_{j}(p_{j})$ are bounded. Therefore, $\mu_{j}(j\in\mathcal{R}_{b})\cdot\phi_{j}(p_{j})$ is Lipschitz in RSS with a Lipschitz constant
$\left(2c_{\phi}+L_{\phi}\right)\cdot\frac{1}{n-d}.$
Moreover, $\sum_{b}T_{b}=\sum_{j=1}^{d}\mu_{j}(j\in\mathcal{R}_{b})\cdot\phi_{j}(p_{j})$ is then Lipschitz in RSS with a Lipschitz constant $Bm\left(2c_{\phi}+L_{\phi}\right)/(n-d)$. A key observation is that ${\sum_{b}T_{b}}/{R}$ is Lipschitz in RSS since $R$ is not a function of RSS.
Now consider the event $A_{R}=\left\{R>Bq\right\}$. Since RSS is independent of $\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$, $A_{\textnormal{RSS}}$ and $A_{R}$ are independent. In addition, we have $\mathbb{P}(A_{R}^{\mathsf{c}})\leq\rho$.
On $A_{\textnormal{RSS}}\cap A_{R}$, we have ${\sum_{b}T_{b}}/{R}$ is $L$-Lipschitz in RSS with
$L=\left(\frac{2mc_{\phi}+mL_{\phi}}{q}\right)\cdot\frac{1}{n-d}.$
@@
$\mathbb{E}\left[\frac{\sum_{b}T_{b}}{R}\mid\textnormal{RSS},A_{R}\right]$
@@
$E_{R}(\textnormal{RSS})=\mathbb{E}\left[\frac{\sum_{b}T_{b}}{R}\mid\textnormal{RSS},A_{R}\right],\quad E_{R^{\mathsf{c}}}(\textnormal{RSS})=\mathbb{E}\left[\frac{\sum_{b}T_{b}}{R}\mid\textnormal{RSS},A_{R}^{\mathsf{c}}\right],$
@@
$E(\textnormal{RSS})=\mathbb{E}\left[\frac{\sum_{b}T_{b}}{R}\mid\textnormal{RSS}\right]=E_{R}(\textnormal{RSS})\cdot\mathbb{P}(A_{R})+E_{R^{\mathsf{c}}}(\textnormal{RSS})\cdot\mathbb{P}(A_{R}^{\mathsf{c}}).$
Note all $E_{R}(\textnormal{RSS})$, $E_{R^{\mathsf{c}}}(\textnormal{RSS})$, and $E(\textnormal{RSS})$ are bounded by $Bc_{\phi}m$.
@@
$=$ $(E_{R}(\textnormal{RSS})-E_{R}(\textnormal{RSS}^{\prime}))^{2}+O(c_{\phi}^{2}m^{2}\cdot B^{2}\rho)$
@@
$\quad+O(c_{\phi}^{2}m^{2}\cdot B^{2}\rho)+O\left(c_{\phi}^{2}m^{2}\cdot B^{2}e^{-{(n-d)}/{32}}\right)$
@@
$=$ $\left(\frac{2mc_{\phi}+mL_{\phi}}{q}\right)^{2}\cdot\frac{2}{n-d},$
where the last line uses the fact that $\textnormal{RSS}\sim\chi^{2}(n-d)$ and hence $\textnormal{Var}(\textnormal{RSS})=2(n-d)$. Then we have
$\textnormal{Var}\left[E(\textnormal{RSS})\right]\leq\left(\frac{2mc_{\phi}+mL_{\phi}}{q}\right)^{2}\cdot\frac{2}{n-d}+O(c_{\phi}^{2}m^{2}\cdot B^{2}\rho)+O\left(c_{\phi}^{2}m^{2}\cdot B^{2}e^{-(n-d)/32}\right).$
@@
$\textnormal{Var}\left[\mathbb{E}\left(\frac{\sum_{b}T_{b}}{R}\mid\textnormal{RSS}\right)\right]\leq\left(\frac{2mc_{\phi}+mL_{\phi}}{q}\right)^{2}\cdot\frac{2}{n-d}+O(c_{\phi}^{2}m^{2}\cdot B^{2}(\rho\vee e^{-(n-d)/32})).$
@@
$\textnormal{Var}\left[\frac{\sum_{b}T_{b}}{R}\right]$ $=\mathbb{E}\left[\textnormal{Var}_{\textnormal{RSS}}\left(\frac{T}{R}\right)\right]+\textnormal{Var}\left[\mathbb{E}\left(\frac{\sum_{b}T_{b}}{R}\mid\textnormal{RSS}\right)\right]$
$\leq O\left(\frac{c_{\phi}^{2}m^{4}}{Bq^{4}}\right)+\left(\frac{2mc_{\phi}+mL_{\phi}}{q}\right)^{2}\cdot\frac{2}{n-d}+O(c_{\phi}^{2}m^{2}\cdot B^{2}(\rho\vee e^{-(n-d)/32}))$
@@
$\textnormal{Var}\left[\widehat{\textnormal{FDR}}\right$]= $\textnormal{Var}\left[\frac{\sum_{b}T_{b}}{\sum_{b}R_{b}}+O\left(\frac{Bc_{\phi}m^{3}}{R^{2}}\right)\right]$
$\leq$ $2\textnormal{Var}\left[\frac{\sum_{b}T_{b}}{\sum_{b}R_{b}}\right]+O\left(2\mathbb{E}\left[\frac{B^{2}c_{\phi}^{2}m^{6}}{R^{4}}\right]\right)$
$\leq$ $2\textnormal{Var}\left[\frac{\sum_{b}T_{b}}{\sum_{b}R_{b}}\right]+O\left(\frac{c_{\phi}^{2}m^{6}}{B^{2}q^{4}}\right)+O(B^{2}c_{\phi}^{2}m^{6}\rho)$
$\leq$ $O\left(c_{\phi}^{2}\left(\frac{m^{4}}{q^{4}}+\frac{m^{6}}{Bq^{4}}\right)\cdot\frac{1}{B}\right)+\left(\frac{2mc_{\phi}+mL_{\phi}}{q}\right)^{2}\cdot\frac{4}{n-d}$
$+O(c_{\phi}^{2}m^{6}\cdot B^{2}(\rho\vee e^{-(n-d)/32})).$
@@
Result (1): The setting satisfies the conditions in Part 2 of Proposition 3.1 with $\xi=B^{-\varepsilon}$ and $q=\xi/3=O(B^{-\varepsilon})$. Then $\rho=\exp\left(-B^{1-\varepsilon}/36\right)$. Plugging them into the bound in Theorem 3.1, we have
$\textnormal{Var}(\widehat{\textnormal{FDR}})=O\left(B^{-1+4\varepsilon}+\frac{B^{2\varepsilon}}{n-d}+B^{2}\exp\left(-B^{1-\varepsilon}/36\right)\right)=O(B^{-1+4\varepsilon}).$